# Neuron Empirical Gradient: Discovering and Quantifying Neurons’ Global Linear Controllability

### v2

@@ I.2 More Data about Efficiency
We report the accuracies of major-vote probers with different neuron sizes for all tasks to provide additional evidence for the discussion about the representation and acquisition efficiency of skill neurons in § H.1. The results are demonstrated in Figure 16 and Figure 17 for zero-shot and few-shot prompting settings.
@@
I.3 Probing With Varying Neuron Sets
We report the aggregated accuracies across all 22 tasks in MCEval8K in Figure 18 to provide additional evidence for discussion in § H.3. It demonstrates that many neurons can construct the classifiers in solving the language tasks, showing their ability to represent language skills and knowledge.
Appendix J Prompting Setups
@@ I.3 Probing With Varying Neuron Sets
Appendix K Diverse Contexts for Skill Neuron Generality Evaluation
@@
We fix the number of skill neurons to 32 when training the skill-neuron-based probers. We use 32 as the optimal neuron size of PAWS with the few-shot setting is 32. Finally, we report the pair-wise generality values among different prompting settings in Figure 19.

### v3

@@ H.1 More Data about Efficiency
We report the accuracies of majority-vote probes with different neuron sizes for all tasks to provide additional evidence for the discussion about the representation and acquisition efficiency of skill neurons in § 6.1. The results are demonstrated in Figure 18 and Figure 19 for zero-shot and few-shot prompting settings.
@@
H.2 Probing With Varying Neuron Sets
We report the aggregated accuracies across all 22 tasks in MCEval8K in Figure 20 to provide additional evidence for discussion in § 6.3. It demonstrates that many neurons can construct the classifiers in solving the language tasks, showing their ability to represent language skills and knowledge.
H.3 Tree-Probe: Flatteness vs. Hierarchy
To investigate the balance between hierarchy and independence of skill neurons, we train Tree-Probe with fixed neuron features ($2^{10}$) but with different depths and trees. For each task, we train 10 Tree-Probes, varying the number of trees (#n_tree $\in(2^{0}\sim 2^{9})$) and the tree depth (#n_layer $\in(2^{10}\sim 2^{1})$), which fewer trees with deeper layers indicate a more hierarchical structure. All tasks show a camel curve given stronger hierarchies. We report the optimal #n_layer for different tasks as follows: CSQA(4), MNLI(16), SWAG(16), Stereoset(16), Agnews(32), MyriadLAMA(32), mLAMA(32), XNLI(32), POS(64), FEVER(64), Toxic(64), LTI(64), GED(128), IMDB(128), M-Amazon(128), CHUNK(256), NER(256), Amazon(256), PAWS(256), HaluEval(256), M-POS(256), TempLAMA(1024). This demonstrates that Different language skills require different hierarchy levels. For instance, factual tasks benefit from flatter structures, while linguistic tasks prefer deeper hierarchies.
For all tasks in MCEval8K, we plot the accuracies of trained models with varying hyperparameters, including the number of trees and layers per tree. The number of trees is set to $2^{N}$, where $N$ ranges from 0 to 10, and the number of layers is set to $2^{M}$, where $M$ ranges from 1 to 11. Training is conducted only for configurations where $N+M<12$. The results are visualized as 3D surfaces, where the x-axis represents the logarithm of the number of trees (#log_ntree), the y-axis shows the logarithm of the number of layers (#log_nlayer), and the z-axis indicates the accuracy evaluated on the test set. We display the results for all tasks under the zero-shot setting in Figure 22 and those under the few-shot setting in Figure 23.
Figure 22: Accuracies of trained random forest models with the zero-shot setting on Llama2-7B.
Figure 23: Accuracies of trained random forest models with the few-shot setting on Llama2-7B.
Appendix I Prompting Setups
@@
Appendix J Diverse Contexts for Skill Neuron Generality Evaluation
@@
We fix the number of skill neurons to 32 when training the skill-neuron-based probes. We use 32 as the optimal neuron size of PAWS with the few-shot setting is 32. Finally, we report the pair-wise generality values among different prompting settings in Figure 21.